You have an order for furosemide 120 mg to be given IM in three equally divided doses every 8 hours. The solution available delivers 20 mg/mL. How many mL would you administer every in each dose?
Answer #1

prescribed dose= 120 mg in 3 divided dose

so ,1 dose = 120 milligram/ 3= 40 milligram= 40 mg

available dose=20 mg/ml

1 millilitre = 20 milligram

prescribed dose in ml for 1 time=

ratio method

20 mg in 1 ml= 40 mg in 'x 'ml

20/1=40/x

x = 40/20 =2 ml

prescribed millilitre per dose = 2 millilitre= 2ml